Property valued at $90,000 is assessed at 2/3 of its value. If the property tax is $30 per $1000 of assessed value per year, what is the yearly property tax bill?
Answer by drk(1908)

Assessed value = 90,000 x (2/3) = 60,000
property tax is
30/1000 = x / 60000
by cross multiplying, we get
1000x = 1800000
x = $1,800.